# TreeCat: Standalone Catalog Engine for Large Data Systems

### v1

@@ 6.1. Discussion
1:  for $(path,predicate)\in txn.ScanSet$ do
@@
3:     if $VersionMap[path]>txn.read\_vid$ then
@@
5:        $iter.seek(path,txn.read\_vid)$
6:        while $iter.path.parent=path$ do
7:           if $iter.type=merge$ and $iter.image.empty()$ then
@@
9:           end if
10:           // precision locking
11:           if $iter.image.before\models predicate$ then
12:              $abort(txn)$
13:           else if $iter.image.after\models predicate$ then
14:              $abort(txn)$
15:           end if
16:        end while
17:     end if
18:  end for
6.2. Implementation
The concurrency control is handled by the TransactionManager, which keeps track of $read\_vid$, the latest version up to which, the catalog can be read safely, and $commit\_vid$, which is the monotone increasing vid assigned to a transaction before final commit. Key in-memory data structures include:
  1. (1)
    $ScanSet$: Maintained by each transaction. Stores pairs of the scan range (in the form of the context object path) and the predicate of the predicate read operations executed as part of the transaction.
  2. (2)
    $LogBuffer$: Maintained by each transaction. Stores before- and after-images of all recent write operations.
  3. (3)
    $VersionMap$: Used by TransactionManager for validation. Maps object path to the most recent $vid$ of its children.
  4. (4)
    $LogIndex$: Used by TransactionManager for validation. Index on the $LogBuffer$, mapping pairs of object path and $vid$ to the log records. The index is implemented using a lock-free skip list and is sorted by the parent path, then by the $vid$.
When $startTransaction()$ is called, a new transaction session is created and $read\_vid$ is loaded as the vid at which all read queries are executed in the transaction session. During query execution, pairs of predicates and parent paths (prefix) of objects scanned for predicate evaluation are added to the $ScanSet$. Also, a private copy of the write set is constructed on the client side. When the transaction is submitted for the commit process, the write set is sent to the server and undergoes preprocessing phase, where it is copied to the $LogBuffer$ and preconditions are checked. If the transaction passes these checks, the transaction is scheduled for the main commit process.
The main commit process consists of two phases: validation and batch write. For each phase, there is a dedicated thread pool of a fixed size, with each thread handling a single hash partition of the $ScanSet$ and the write set. At the start of the validation, TransactionManager executes a $fetch\_and\_add()$ operation on the $commit\_vid$ and assigns it to the transaction. Then, each worker thread validates the scan ranges in the $ScanSet$ against the $VersionMap$ to determine whether any objects scanned by predicate read operations have changed. If potential conflicts are detected, an iterator on $LogIndex$ is set to scan and apply precision locks on corresponding recent writes with $vid$ greater than the transaction’s $read\_vid$ and fall in the scan range. If any conflict is detected through precision locking, the transaction is aborted. After all threads pass validation, they update the $VersionMap$ and $LogIndex$ based on their respective write set partitions for the transaction. The validation phase (omitting the initial $vid$ assignment and latter updates to in-memory data structures) proceeds as per Algorithm  1. Note that for merge operation, before- and after-image may have to be constructed as they are not available until the final commit. However, once constructed, the images are cached so they are accessible during the validation of subsequent transactions.
Once validation is complete, the transaction proceeds to the write phase. During this phase, each thread processes its assigned write partition, evaluating the final values of merge operations (commit-time updates) and applying the writes as an atomic batch write operation to the underlying RocksDB instance. After every thread finishes its batch write, WAL records of the transaction are flushed to disk along with those of other transactions via $SyncWAL()$ operation, which calls $fsync()$ on the WAL files. The $read\_vid$ is also updated, making the committed changes visible for subsequent reads. Transactions are processed one at a time in each phase, but parallelism is achieved via hash-partitioning the workload and pipelining different phases of the commit process.
Old $LogBuffer$s and entries in $VersionMap$ and $LogIndex$ that are no longer needed are asynchronously garbage collected by a background thread. For this purpose, the TransactionManager keeps track of $watermark$, the oldest $read\_vid$ among the current transactions. After each transaction commit, the $watermark$ is updated and all entries in the aforementioned data structures with $vid$ below the $watermark$ are garbage collected.
@@ 6.3. Proof of Correctness
Concurrency Control Model: We first formalize the concurrency control model. We follow Adya’s model (Adya et al., 2000), where every object in the database has one or more versions. Every operation on an object is mapped to exactly one version of that object. Given an object $x$, $x_{i}$ is a version with $vid$ $i$, installed by transaction $T_{i}$. Every object is assumed to exist forever regardless of when they are inserted or deleted. The unborn state is represented as version $x_{init}$, while the dead state is represented as $x_{i}$, installed by the corresponding delete operation of $T_{i}$. Every object is uniquely identified by its full path.
It is important to discuss how TreeCat’s query evaluation is represented in this model. Evaluation of a path expression involves performing a series of predicate range scans, each with respect to a context parent object. For example, given the path expression, /[obj_id = retail]/[name = "Sales"], the evaluation of the second predicate, $name=``Sales"$, is executed as a range scan on the children of /retail. Each range scan can be modeled as a predicate read operation, where the predicate is a conjunction of the original predicate and the equivalence between the object parent path and the context parent path. Following the above example, evaluating $name=``Sales"$ against /retail’s children objects can be modeled as a predicate read operation with the predicate, $name=``Sales"\wedge parent=/retail$. It should be noted that the evaluation of a single predicate in a path query can correspond to multiple such predicate read operations, as there can be multiple context parent objects with which the predicate is evaluated.
@@
We prove the lemma by enumeration. First, there cannot be any ww-dependency from $T_{j}$ to $T_{i}$ because write operations are performed in the order of the $commit\_vid$ in the write phase. Second, wr-dependency from $T_{j}$ to $T_{i}$ is also not possible. Given that $read\_vid_{i}<commit\_vid_{i}$ and $commit\_vid_{i}<commit\_vid_{j}$, $read\_vid_{i}<commit\_vid_{j}$ by transitivity. Because changes committed by $T_{j}$ become visible only after the $read\_vid$ has been incremented to $commit\_vid_{j}$, $T_{i}$, which has a lower $read\_vid$, cannot observe any changes installed by $T_{j}$. Lastly, we show by contradiction that there cannot be any rw-antidependency from $T_{j}$ to $T_{i}$. Assume that there is a rw-antidependency from $T_{j}$ to $T_{i}$. Consider the state when $T_{i}$ passed validation, but $T_{j}$ is about to enter validation. There must be a write operation $w$ executed by $T_{i}$ on an object $o$, whose before- or after-image satisfies the predicate of a predicate read operation $P$ executed by $T_{j}$, which scans $w$’s before-image. Per our earlier discussion, the predicate of $P$ is $p\wedge parent=c$, where $p$ is the predicate in the path query and $c$ is the context parent object. During the query execution of $T_{j}$, $(c,p)$ was added to $T_{j}$’s $ScanSet$. Also, $(o.parent,i)$ was added to $VersionMap$ and a mapping from $o$ to $w$’s before- and after-images was added to $LogIndex$ during the validation of $T_{i}$. During the validation of $T_{j}$, $c$ is found in $VersionMap$ because $c=o.parent$ holds true. Since $read\_vid_{j}<commit\_vid_{i}$, $p$ is evaluated against $o$’s entries in $LogIndex$. Because the before- or after-image of $w$ satisfies $p$, the rw-conflict is detected and $T_{j}$ is aborted, leading to a contradiction. We have now considered all types of dependencies, so the lemma holds true. ∎
@@
We prove the theorem by contradiction. Assume that there is a schedule $s$, output by TreeCat, that is not conflict serializable. Then, the corresponding precedence graph, $G(s)=(V,E)$ must contain at least one cycle. Consider a directed edge $(T_{i},T_{j})$ in the cycle. Because there is a dependency from $T_{i}$ to $T_{j}$, $commit\_vid_{i}<commit\_vid_{j}$ by Lemma 1. Because $(T_{i},T_{j})$ is in a cycle there is a path from $T_{j}$ to $T_{i}$. Using Lemma 1 and transitivity, we can prove by induction (which we omit for brevity) that $commit\_vid_{i}>commit\_vid_{j}$. We end up with a contradiction, so the theorem holds true. ∎
System MVCC Optimistic Pessimistic Multi-table Transaction
AWS Redshift
Google BigQuery
Snowflake
Vertica
HMS
Delta Lake
Iceberg Nessie Catalog
Hudi
TreeCat
Table 1. Concurrency control mechanisms of OLAP systems.
@@ 6.4. Related Work
Catalog Concurrency Control: Traditional RDBMSs store catalog as a set of relations, relying on the same concurrency mechanism that is used for data. While data warehouses use proprietary mechanisms, Snowflake is known to use FoundationDB to persist transaction states and lock queues (Motivala, 2018). Even though resource locks at the level of micro-partitions are available, most single table DML queries require locking all micro-partitions (Snowflake, 2023). Delta Lake uses snapshot-based OCC where, during the final commit, it tries to atomically write the next delta log record. If the underlying storage system does not support atomic file rename, it has to resort to another transactional system, such as DynamoDB (Scott Sandre and Kryński, 2022). Delta Lake currently only supports single-table transactions. However, there is early work on introducing a new commit protocol called commit coordinator, an external system that is dedicated to handling concurrency control, which may support multi-table transactions in the future (Tathagata Das et al., 2024). Iceberg also uses snapshot-based OCC where, after writing all the metadata files (manifest files, manifest list files, and metadata file), it does a final commit by switching the current pointer in the catalog to the new metadata file. If there is a conflict, the transaction undergoes a retry validation, so the entire operation need not be repeated from scratch (Okolnychyi et al., 2024). Like Delta Lake, Iceberg’s commit protocol only supports single-table transactions. Recently, multi-table transactions have been added to the Iceberg REST API. But as far as we know, Nessie (Nessie, [n.d.]), which supports git-like versioning, is the only compatible catalog that supports it. Even so, the transactions operate at the level of a single branch of the whole database. Lastly, Hudi also uses snapshot-based OCC and does conflict resolution based on the files modified by multiple writers (Foundation, 2024b). Hudi requires an external lock provider for locking tables and can only support single-table transactions.
@@
Predicate Read: In order for a data system that supports predicate read operations to guarantee serializable isolation, it has to account for any predicate dependencies in addition to item dependencies (Adya et al., 2000). The standard solution is to lock the entire scan range of the predicate read operation, an example of which is next key locking (Mohan et al., 1989), which locks the current record and the gap until the next record. Since taking many record locks can cause excessive overheads, multiple granularity locking (Gray et al., 1976) is commonly used to escalate locks to more coarse-grained ones, such as table locks. Large scan range may cause unnecessary conflicts, so it is crucial to minimize the scan range. For simple predicates, performing an index scan can limit the scan range. A more advanced technique may be used to better approximate the minimum scan range for complex predicates. For example, Spanner uses a runtime data structure called a filter tree to compute the approximate scan range (Bacon et al., 2017). One way to precisely handle predicate dependencies is to validate the scan operation by re-executing it, which is adopted by Hekaton (Diaconu et al., 2013). Another approach is to use predicate-based locking. Conventional predicate locking involves satisfiability test problem, which is known to be NP-complete (Hunt and Rosenkrantz, 1979). However, precision locking avoids this problem by directly evaluating predicate against before- and after-images of posted writes (Jordan et al., 1981). The technique was not used in any real system for a long time, but was later adopted by Neumann et al. for implementing the MVOCC mechanism of Hyper (Neumann et al., 2015).
@@ 7.1. Concurrency Control Mechanism
Figure 9. Throughput for read-intensive (80-20) and write-intensive (20-80) workloads with multiple databases (10 client threads per database), respectively.
Lastly, we evaluate how throughput scales across different numbers of databases for two types of workload with low contention (10 threads per database). As expected, all three schemes have very similar throughput for the read-intensive workload, saturating around 3 databases. For the write-intensive workload, OSPL generally outperforms the others. However, the throughput converges with that of MGL as it is increasingly constrained by contention on physical resources. OSPL still outperforms MGL at 6 databases by a small margin. OSL exhibits significantly lower throughput, compared to the other schemes, and also reaches the saturation point earlier at 3 databases due to high abort rate.
@@ 7.2. Comparative Evaluation
Next, we compare TreeCat with HMS, Delta Lake, and Iceberg. An apples-to-apples comparison is difficult because TreeCat does not support a catalog API fully compatible with Spark and all the systems have different architectures and transactional guarantees. We nonetheless try our best to compare the relative costs of metadata operations. We set up HDFS (Hadoop Distributed File System) and Spark on the 4 node cluster. We used Hive 2.3, Delta Lake 2.4, and Iceberg 1.5, all of which are compatible with Spark 3.4. We used a single table with a schema identical to that of the Store Sales table in the TPC-DS, partitioned by the sales date, from 1998-01-01 to 2003-12-31. To test scalability of read performance, we populated the table with varying numbers of files and measured the latency of retrieving files with a range filter on the date. While our primary implementation of the catalog API uses a single thread to retrieve the query result via a single stream, both HMS and Iceberg leverage multiple cores on the client machine to retrieve the result over multiple streams where HMS calls $listFile()$ and Iceberg retrieves manifest files in parallel. Client-side parallelism can be achieved with TreeCat quite easily by adding a partitioning predicate to the path query, so that multiple threads can concurrently submit queries and retrieve distinct result sets. Although a hash function combined with the $mod$ operation would typically be the preferred approach, our implementation applies the $endwith()$ operation to the partition value due to the limitations of the current query language. We used this parallel implementation for the following experiments.
@@
We first analyze the scalability of the file retrieval operation with high selectivity (date range of 1 day) and with low selectivity (date range of 365) across different numbers of files. For the 1-day predicate range, TreeCat consistently shows low latency as it efficiently traverses the hierarchical data on the server and only returns the file metadata in the given data partition. HMS also scales quite well as it calls $listPartitionsbyFilter()$ to only retrieve the partitions in the given range from the thrift server. Although Delta Lake benefits from Spark’s distributed processing engine, the overhead of setting up and executing the job adds significant latency, resulting in a relatively high minimum latency of around 600 ms. This contrasts with previous experimental results (Armbrust et al., 2020; Jain et al., 2023) and we suspect that performance could be significantly improved with a better-configured, larger Spark cluster. But the client-side implementation of getting the end result appears to be a major bottleneck. Although Iceberg utilizes multiple cores for processing manifest files, the size of an individual manifest file can get large, which contributes to the longer latency for the higher file counts. TreeCat outperforms the other systems and scales well also for the 365-days predicate range. HMS beats both Delta Lake and Iceberg when the file count reaches 50000, which we mainly attribute to the relatively low latency of HDFS, compared to object stores.
Figure 11. Median latency of file retrieval across different predicate ranges, and numbers of physical cores on the client machine, respectively. File count is 50000 for both while predicate range is 365 for the latter.
@@
Figure 12. Throughput and latencies of $alterTable()$.
Lastly, we measure the relative cost of the commit protocol for each system by invoking a light-weight metadata operation, $alterTable()$, on an empty Store Sales table. The experiment is only intended to provide a rough estimate as the systems have different client-side implementations and provide different transactional guarantees (which HMS does not even provide). We measure both the total throughput over a set duration and end-to-end latency of each invocation. We observe that lakehouse storage formats, especially Delta Lake, have significantly longer latencies with median latency of 1434.1 ms for Delta Lake and 180.4 ms for Iceberg (throughput of 0.6 ops/s and 5.6 ops/s, respectively). Delta Lake and Iceberg first have to reconstruct the most current snapshot of the table, which tends to be costly. In addition, the final commit protocol involves creating metadata files and calling an atomic operation on the underlying HDFS, which contributes to the extra latency. As a transactional system, TreeCat exhibits significantly lower median latency of 8.4 ms and higher throughput of 117 ops/s.
8. Conclusion
Motivated by the increasing use of a diverse set of data management systems within a single enterprise and the rise of disaggregated architectures, in this paper, we identified the critical systemic and functional requirements for a standalone catalog and justified the key design decisions for our catalog engine, TreeCat. By adopting a hierarchical data model, introducing a specialized data format and layout, and developing a novel concurrency control mechanism, we directly address emerging use cases that challenge state-of-the-art operational metadata systems. Our experimental evaluation not only exposes the limitations of existing solutions but also validates the advantages of our design choices. Looking ahead, we are actively pursuing further enhancements, including (1) addressing additional requirements outlined in Section 3, (2) implementing and testing more complex workloads—such as processing non-tabular data, provenance, and versioning, and (3) developing a comprehensive catalog API to enable complete end-to-end evaluation.

### v2

@@ 5.2. Implementation
1: for $(path,predicate)\in txn.ScanSet$
@@
3:     if VersionMap[path]>txn.read_vidVersionMap[path]>txn.read\_vid
@@
5:        iter.seek(path,txn.read_vid)iter.seek(path,txn.read\_vid)
6:        while $iter.path.parent=path$
7:          if $iter.type=merge$ and $iter.image.empty()$
@@
9:          // precision locking
10:          if $iter.image.before\models predicate$
11:            $abort(txn)$
12:          else if $iter.image.after\models predicate$
13:            $abort(txn)$
@@ 5.3. Proof of Correctness
Concurrency Control Model: We first formalize the concurrency control model. We follow Adya’s model (Adya et al., 2000), where every object in the database has one or more versions. Every operation on an object is mapped to exactly one version of that object. Given an object $x$, $x_{i}$ is a version with vidvid $i$, installed by transaction $T_{i}$. Every object is assumed to exist forever regardless of when they are inserted or deleted. The unborn state is represented as version $x_{init}$, while the dead state is represented as $x_{i}$, installed by the delete operation of $T_{i}$. Every object is uniquely identified by its full path.
It is important to discuss how TreeCat’s query evaluation is represented in this model. Evaluation of a path expression involves performing a series of predicate range scans, each with respect to a context parent object. For example, given the path expression, /[obj_id = ‘‘retail"]/[name = ‘‘sales"], the evaluation of the second predicate, name=salesname=``sales", is executed as a range scan on the children of /retail. Each range scan can be modeled as a predicate read operation, where the predicate is a conjunction of the original predicate and the equivalence between the object parent path and the context parent path. Following the above example, evaluating name=salesname=``sales" against /retail’s children objects can be modeled as a predicate read operation with the predicate, $name=``sales"\wedge parent=/retail$. The evaluation of a single predicate in a path query can correspond to multiple such predicate read operations, as there can be multiple context parent objects with which it is evaluated.
@@
We prove the lemma by enumeration. First, there cannot be any ww-dependency from $T_{j}$ to $T_{i}$ because write operations are performed in the order of the commit_vidcommit\_vid in the write phase. Second, wr-dependency from $T_{j}$ to $T_{i}$ is also not possible. Given that $read\_vid_{i}<commit\_vid_{i}$ and $commit\_vid_{i}<commit\_vid_{j}$, $read\_vid_{i}<commit\_vid_{j}$ by transitivity. Because changes committed by $T_{j}$ become visible only after the read_vidread\_vid has been incremented to $commit\_vid_{j}$, $T_{i}$, which has a lower read_vidread\_vid, cannot observe any changes installed by $T_{j}$. Lastly, we show by contradiction that there cannot be any rw-antidependency from $T_{j}$ to $T_{i}$. Assume that there is a rw-antidependency from $T_{j}$ to $T_{i}$. Consider the state when $T_{i}$ passed validation, but $T_{j}$ is about to enter validation. There must be a write operation $w$ executed by $T_{i}$ on an object $o$, whose before- or after-image satisfies the predicate of a predicate read operation $P$ executed by $T_{j}$, which scans $w$’s before-image. Per our earlier discussion, the predicate of $P$ is $p\wedge parent=c$, where $p$ is the predicate in the path query and $c$ is the context parent object. During the query execution of $T_{j}$, $(c,p)$ was added to $T_{j}$’s $ScanSet$. Also, $(o.parent,i)$ was added to $VersionMap$ and a mapping from $o$ to $w$’s before- and after-images was added to $LogIndex$ during the validation of $T_{i}$. During the validation of $T_{j}$, $c$ is found in $VersionMap$ because $c=o.parent$ holds true. Since $read\_vid_{j}<commit\_vid_{i}$, $p$ is evaluated against $o$’s entries in $LogIndex$. Because the before- or after-image of $w$ satisfies $p$, the rw-conflict is detected and $T_{j}$ is aborted, leading to a contradiction. We have now considered all types of dependencies, so the lemma holds true. ∎
@@
We prove the theorem by contradiction. Assume that there is a schedule $s$, output by TreeCat, that is not conflict serializable. Then, the corresponding precedence graph, $G(s)=(V,E)$ must contain at least one cycle. Consider a directed edge $(T_{i},T_{j})$ in the cycle. Because there is a dependency from $T_{i}$ to $T_{j}$, $commit\_vid_{i}<commit\_vid_{j}$ by Lemma 1. Because $(T_{i},T_{j})$ is in a cycle there is a path from $T_{j}$ to $T_{i}$. Using Lemma 1 and transitivity, we can prove by induction (which we omit for brevity) that $commit\_vid_{i}>commit\_vid_{j}$, resulting in a contradiction. ∎
@@ 5.4. Related Work
Catalog Concurrency Control: Traditional RDBMSs store the catalog as a set of relations, relying on the same concurrency mechanism that is used for data. While data warehouses use proprietary mechanisms, Snowflake is known to use FoundationDB to persist transaction states and lock queues (Motivala, 2018). Even though resource locks at the level of micro-partitions are available, most single table DML queries require locking all micro-partitions (Snowflake, 2023). Delta Lake uses snapshot-based OCC where, during the final commit, it tries to atomically write the next delta log record. If the underlying storage system does not support atomic file rename, it has to resort to another transactional system, such as DynamoDB (Scott Sandre and Kryński, 2022). Delta Lake currently only supports single-table transactions. However, there is early work on introducing a new commit protocol called commit coordinator, an external system that is dedicated to handling concurrency control, which may support multi-table transactions in the future (Tathagata Das et al., 2024). Iceberg also uses snapshot-based OCC where, after writing all the metadata files (manifest files, manifest list files, and metadata file), it does a final commit by switching the current pointer in the catalog to the new metadata file. If there is a conflict, the transaction undergoes a retry validation, so the entire operation need not be repeated from scratch (Okolnychyi et al., 2024). Like Delta Lake, Iceberg’s commit protocol only supports single-table transactions. Recently, multi-table transactions have been added to the Iceberg REST API. But as far as we know, Nessie (Nessie, 2025), which supports git-like versioning, is the only compatible catalog that supports it. Even so, the transactions operate at the level of a single branch of the whole database. Lastly, Hudi also uses snapshot-based OCC and does conflict resolution based on the files modified by multiple writers (Foundation, 2025b). Hudi requires an external lock provider for locking tables and can only support single-table transactions.
@@
Predicate Read: To guarantee serializable isolation, a data system must account for any predicate dependencies (Adya et al., 2000). The standard solution is to lock the entire scan range of the predicate read operation, an example of which is next key locking (Mohan, 1990), which locks the current record and the gap until the next record. Since taking many record locks can cause excessive overheads, multiple granularity locking (Gray et al., 1994) is commonly used to escalate locks to more coarse-grained ones, such as table locks. Large scan range may cause unnecessary conflicts, so it is crucial to minimize the scan range. For simple predicates, performing an index scan can limit the scan range. A more advanced technique may be used to better approximate the minimum scan range for complex predicates. For example, Spanner uses a runtime data structure called a filter tree to compute the approximate scan range (Bacon et al., 2017). One way to precisely handle predicate dependencies is to validate the scan operation by re-executing it, which is adopted by Hekaton (Diaconu et al., 2013). Another approach is to use predicate-based locking. Conventional predicate locking involves satisfiability test problem, which is known to be NP-complete (Hunt and Rosenkrantz, 1979). However, precision locking avoids this problem by directly evaluating predicate against before- and after-images of posted writes (Jordan et al., 1981). The technique was not used in any real system for a long time, but was later adopted by Neumann et al. for implementing the MVOCC mechanism of Hyper (Neumann et al., 2015).
@@ 6.1. Concurrency Control Mechanism
Lastly, we evaluate how two important structural properties of the metadata schema, breadth and depth, affect performance. These experiments use a separate microbenchmark designed to be more generalizable. The workload consists of a mix of read and read-write operations (50-50) on a single data set, which is range-partitioned into 100000 data files by a set of independent clustering attributes. Both types of operation perform a predicate read operation with fixed selectivity, using range predicates on randomly selected clustering attributes. Read-write operation additionally inserts a new data file. A partition level is constructed by grouping data files by a clustering attribute. Nested grouping by multiple attributes results in a multi-level hierarchy where the number of such attributes determines the overall partition level (depth) and the cardinality determines the fan-out (breadth). As shown in Figure 9, throughput is low across all schemes at low fan-out because of large scan ranges (e.g., all 100000 files have to be scanned at the partition level of 1). However, both OSL and MGL are further affected by contention from coarse-grained scan range locking. OSPL, on the other hand, avoids false conflicts as it uses a predicate-based method. Throughput converges with higher fan-out, as MGL and OSL benefit from more fine-grained lock ranges. A similar trend can be observed with different partition levels (depths) for the same reason. However, throughput is divergent even at the highest partition level because a subset of the clustering attributes do not form partitioning groups; a predicate read operation only on these attributes requires scanning all partition groups, resulting in higher contention for OSL and MGL. Although OSL outperforms MGL in throughput, it suffers from high abort rates.
Figure 9. Throughput for different breadths (fan-out) and depths (partition level) of the metadata schema.
@@ 6.2. Comparative Evaluation
Next, we compare TreeCat with HMS, Delta Lake, and Iceberg. An apples-to-apples comparison is difficult because TreeCat does not support a catalog API fully compatible with Spark and all the systems have different architectures and transactional guarantees. We nonetheless try our best to compare the relative costs of metadata operations and conduct an end-to-end experiment with an alternate database engine. We set up HDFS (Hadoop Distributed File System) and Spark on the 4 node cluster. We used Hive 2.3, Delta Lake 2.4, and Iceberg 1.5, all of which are compatible with Spark 3.4. We used a single table with a schema identical to that of the Store Sales table in the TPC-DS, partitioned by the sales date, from 1998-01-01 to 2003-12-31. To test scalability of read performance, we populated the table with varying numbers of files and measured the latency of retrieving files with a range filter on the date. While our primary implementation of the catalog API uses a single thread to retrieve the query result via a single stream, both HMS and Iceberg leverage multiple cores on the client machine to retrieve the result over multiple streams where HMS calls $listFile()$ and Iceberg retrieves manifest files in parallel. Client-side parallelism can be achieved with TreeCat quite easily by adding a partitioning predicate to the path query, so that multiple threads can concurrently submit queries and retrieve distinct result sets. Our implementation applies the $endwith()$ operation to the partition value due to the limitations of the current query language. We used this parallel implementation for the following experiments.
@@
As shown in Figure 10, we first analyze the scalability of the file retrieval operation with high selectivity (date range of 1 day) and with low selectivity (date range of 365) across different numbers of files. For the 1-day predicate range, TreeCat consistently shows low latency as it efficiently traverses the hierarchical data on the server and only returns the file metadata in the given data partition. HMS also scales quite well as it calls $listPartitionsbyFilter()$ to only retrieve the partitions in the given range from the thrift server. Although Delta Lake benefits from Spark’s distributed processing engine, the overhead of executing the job adds significant latency, resulting in a relatively high minimum latency of around 600 ms. This contrasts with previous experimental results (Armbrust et al., 2020; Jain et al., 2023) and we suspect that performance could be significantly improved with a better-configured, larger Spark cluster. But the client-side implementation of getting the end result appears to be a major bottleneck. Although Iceberg utilizes multiple cores for processing manifest files, the manifest file size can get large, resulting in the longer latency for the higher file counts. TreeCat outperforms the other systems and scales well also for the 365-days predicate range. HMS beats both Delta Lake and Iceberg when the file count reaches 50000, which we mainly attribute to the low latency of HDFS.
Figure 11. Median latency of file retrieval across different predicate ranges and numbers of physical cores on client machine, resp. File count is 500000 for both while predicate range is 365 for the latter.
@@
We conclude with an end-to-end experiment, using DuckDB 1.2 as the execution engine. The workload involves repeatedly inserting data of constant size into an unpartitioned Store Sales table. A single operation proceeds in three steps: (1) Retrieve the table metadata, including the table directory and schema from the catalog; (2) Parse the schema, generate data, and write the output to a parquet file under the table directory, using DuckDB; (3) Invoke the commit protocol on the catalog with the metadata of the new data file. We measure both the total throughput over a set duration and the end-to-end latency of each invocation across different data file sizes, configured using the number of rows per file (powers of 10, starting from 100). The results are shown in Figure 12. We observe that the high latency of metadata operations can indeed cause a performance bottleneck, especially for smaller file sizes. As shown in the time breakdown of Figure 12, the main difference lies in the commit protocol: Delta Lake’s commit protocol takes over 1.5 seconds, and Iceberg takes around 250 ms, while TreeCat only takes around 10 ms. The commit protocol for lakehouse storage formats involves retrieving the latest table snapshot (which may require reading and processing metadata files), writing one or more metadata files, and invoking a CAS (compare-and-swap) operation for ACID guarantees. The entire process involves one or more operations on the shared storage system, which have non-trivial overhead. Delta Lake incurs extra overhead from launching a Spark job for metadata processing. In contrast, TreeCat uses in-memory data structures for validation, applies writes to RocksDB, which is optimized for fast write, and is only delayed by fsync() operation on the WAL file, resulting in lower latency. The throughput eventually converges with larger file sizes as the cost of the data operation becomes more dominant. However, given applications such as streaming systems (Carbone et al., 2015; Armbrust et al., 2018) that generate small data files at high frequency, these results carry significant real-world implications.
Figure 12. Throughput of insert operations (log scale) for different write sizes, and time breakdown for an insert operation of 10000 row data file ($\approx$ 238 KB), respectively.
7. Conclusion and Discussion
In this paper, we identified the primary use cases of a standalone operational catalog and justified the key design decisions for our catalog engine, TreeCat. Our experimental evaluation not only exposes the limitations of existing solutions but also validates the advantages of our design choices.
Looking ahead, we see a number of critical challenges in this space that need further research. Some of these challenges lie outside the scope of the catalog itself. For instance, consistency between the data and the metadata (e.g., ensuring that a pointer to a physical object is valid), is a major emerging challenge. Yet it is not clear where that responsibility lies. Similarly, access control is crucial but needs to be enforced by the layer that manages access to the data itself. Garbage collection of data objects and their metadata that are no longer visible may become necessary over time. A TreeCat instance can be periodically scanned to find and return leaf objects that are no longer reachable by any path from the root. These objects can be quarantined in temporary storage until the data layer invokes a final commit to remove them. Finally, scalability and availability are important concerns in a distributed setting. Given the relatively small volume of metadata and large numbers of read queries, we propose the adoption of a single-writer, multiple-reader architecture where all read-write transactions are handled by a single server, and logs are shipped to and replayed by read replicas  (Verbitski et al., 2017). This scheme would require minimum changes to the concurrency control mechanism because a single server will process all write operations. If the primary writer server fails, a new writer server can be elected via a quorum-based mechanism, ensuring availability.
Acknowledgements.
This research was supported in part by Dolby Laboratories.